OpenStudy (anonymous):

find real numbers a and b such that the equation is true (a-1)+(b+3)i=5+8i

OpenStudy (sidmanfu):

In equations like this you can always separate real and imaginary numbers into separate equations. So, hypothetically if you had x+yi = 4 - 3i you could separate the equation to x=4 and yi= -3i, so x=4 and y= -3 Now back to your question: First you multiply things out (this isn't that important for the above question, but could be if you had to keep track of multiplying i's together i.e. x + (y+zi)*(i+3) = 8. I won't solve it, but you would have to "FOIL" that last part.) so... back to your question again. multiplying stuff out... a - 1 + b*i +3*i = 5 + 8*i separate into separate equations for real and imaginary (as shown at top) a - 1 = 5 b*i + 3*i = 8*i now solve using basic algebra to get... a = 6 b*i = 5*i, therefore b=5

OpenStudy (anonymous):

No direct answers, please.

OpenStudy (sidmanfu):

I hope I helped enough @Patsfan12 :)

OpenStudy (anonymous):

thanks! your explanation was very thorough :)

OpenStudy (sidmanfu):

no problem! ask me any questions u help with ;)

OpenStudy (anonymous):

i will don't worry XD

OpenStudy (sidmanfu):

lol

OpenStudy (sidmanfu):

and please fan me

OpenStudy (anonymous):

I thought I did?

OpenStudy (anonymous):

yes I am :)

OpenStudy (sidmanfu):

ok thanks :P

OpenStudy (anonymous):

Ok @sidmanfu the instructions are to perform the operation and write the result in standard form: sqrt(-6) x sqrt(-2)

OpenStudy (triciaal):